A team of cancer research specialists received a grant of 300,000 to be used for cancer research. They deposited some of the money in a 10% simple internet account and the remainder in an 8.5% annual simple interest account. How much was deposited in each account if the annual interest is 28,500 ?
Answer by richwmiller(17219)

x+y=300000,
0.1*x+0.085*y=28500
x=300000-y
0.1*(300000-y)+0.085*y=28500
30000-0.1y+0.085*y=28500
-0.015*y=-1500
y=100000
x=300000-y
x=200000
check
0.1*200000+0.085*100000=28500
20000+8500.0=28500
28500.0=28500
ok